The speed of a particle moving along a circular path is decreasing at the rate of 3 m/s^2 .If radius of the circle is 4m,then what'll be the instantaneous acceleration of the particle when it's speed is 4m/s ?


A) 4 towards the center


B) 3 along the tangent


C) 5 by making 37 degrees with tangent


D) 5 by making 53 degrees with the tangent

Grade:12

1 Answers

Chetan Mandayam Nayakar
13 years ago

acceleration=sqrt(32+(42/4)2=5m/s2 at arctan(4/3) =53 degrees with tangent